# TokLIP: Marry Visual Tokens to CLIP for Multimodal Comprehension and Generation

### v1

@@ 4.1 Experimental Setup
The TokLIP architecture integrates the text-to-image VQGAN tokenizer from LlamaGen [47] and a token encoder with ViT-So400M architecture. The downsample ratio is 16. Input resolutions are 256×256 for the base model and 384×384 for the large variant, resulting in a 256 and 576 tokens accordingly. We initialized the weights of token encoder from SigLIP2 [54] as a good basic for semantic perception. A two-layer MLP bridges the VQGAN and CLIP-based token encoder, mapping 8-dimensional VQGAN features as the input of token encoder. The TokLIP-large model trains on 80 million image-text pairs from Capfusion [67] and CC12M [5] and LAION-high-resolution, while the TokLIP-base model augments this with subsets of LAION400M [45], yielding a combined 125 million samples. We filter low-resolution images to ensure effective learning.
For multimodal comprehension, we utilize TokLIP as the input encoder and pass the output tokens through a 2-layer MLP to Qwen2.5-7B-Instruct [64]. Our training process consists of two stages. In Stage 1, we jointly train the MLP and the LLM on 4.5 million images with detailed captions for multimodal alignment, using data sourced from LLaVA-OneVision [26]. In Stage 2, we perform instruction tuning using the LLaVA-v1.5-mix-665K dataset [35]. For visual generation, we follow LlamaGen [47], training class-to-image generation models on ImageNet. We replace LlamaGen’s text-to-image VQGAN tokenizer with TokLIP.
@@
We evaluate the understanding capacity of our TokLIP on ImageNet classification [11] and MSCOCO 5K retrieval [8] tasks. For multimodal comprehension, we rigorously evaluate our model on several zero-shot vision-language benchmarks, which include POPE [29], MME-P (MME-Perception) [17], MMB (MMBench) [36], SEED (SEED-Bench Img) [27], GQA [21], MMMU [69], and MM-Vet [68], covering a diverse range of tasks that test both visual and linguistic understanding in a variety of real-world settings. For visual generation, we mainly report Fréchet inception distance (FID) on the ImageNet 256x256 benchmark.
Table 1: Comparative analysis of tokenizers on multimodal comprehension tasks. We adopt the LLaVA-v1.5 [35] training framework and combine tokenizers with Qwen2.5-7B-Instruct [64], training on LLaVA-v1.5 data. denotes the results are reported from the original paper.
@@ 4.2 Comparison with Other Tokenizers
We compare TokLIP with state-of-the-art tokenizers [61, 71] and CLIP models [9, 43, 70] across both image classification and cross-modal retrieval tasks, shown in Table 2. It can be observed that TokLIP outperforms several SOTA CLIP-based encoders. For instance, TokLIP achieves an impressive 76.4% accuracy on ImageNet with a 256 input resolution, surpassing MetaCLIP [63] and EVA02 [16]. These results demonstrate that TokLIP effectively inherits semantic understanding from the pre-trained teacher and demonstrates competitive performance on downstream tasks. Additionally, at both 256x256 and 384x384 input resolutions, TokLIP outperforms VILA-U [61] and QLIP [71], particularly excelling in the retrieval task. TokLIP achieves 68.00% TR@1 and 52.87% IR@1, which are 7.14% and 9.87% higher than QLIP, respectively, at a 384 resolution. This suggests that TokLIP captures more semantic information compared to methods that discretize high-level image features. Notably, VILA-U is trained on COYO-700M [3], and QLIP uses DataComp-1B [18], both of which involve significantly more resource-intensive training processes than ours. Future work will focus on enhancing TokLIP’s performance through the utilization of higher-quality training data [15, 63].
Model Dataset Res. IN Top1$\uparrow$ COCO TR@1$\uparrow$ COCO IR@1$\uparrow$
EVA02-B/16 [16] Merged-2B 224 74.7 58.74 42.15
CLIP-L/14 [43] WIT400M 224 75.5 56.34 36.51
MetaCLIP-L/14 [63] MetaCLIP400M 224 76.2 60.00 43.81
VILA-U [61] COYO700M 256 73.3 62.56 45.37
QLIP [71] DataComp1B 256 74.3 55.62 38.91
TokLIP Mix 125M 256 76.4 64.06 48.46
OpenCLIP-ViT-L/14 [9] LAION2B 336 75.3 63.36 46.51
SigLIP-B/16 [70] WEBIL10B 384 76.5 67.74 49.90
SigLIP-B/16 [70] WEBIL10B 512 79.1 68.72 50.55
VILA-U [61] COYO700M 384 78.0 - -
QLIP [71] DataComp1B 384 79.1 60.86 43.00
TokLIP Mix 80M 384 80.0 68.00 52.87
@@
LLaVA-v1.51.51.51.5-Phi-$1.5$ [62] $1.3$B 256 84.1 1128.0 - - 56.5 30.7 -
LLaVA-Phi [74] $2.7$B 336 85.0 1335.1 59.8 - - - 28.9
LLaVA-v1.51.51.51.5 [35] $7$B 336 85.9 1510.7 64.3 58.6 62.0 35.4 31.1
InstructBLIP [10] $7$B 224 - - 36.0 53.4 49.2 - 26.2
Qwen-VL-Chat [2] $7$B 448 - 1487.5 60.6 58.2 57.5 - -
IDEFICS-$9$[24] $8$B 224 - - 48.2 - 38.4 - -
InstructBLIP [10] $13$B 224 78.9 1212.8 - - 49.5 - 25.6
@@
DreamLLM [12] $7$B 224 - - - - - - 36.6
LaVIT [22] $7$B 224 - - 58.0 - 46.8 - -
MetaMorph [53] $8$B 384 - - 75.2 71.8 - - -
NExT-GPT [60] $13$B 224 - - - 57.5 - - -
SEED-X [20] $13$B 448 84.1 1457.0 70.1 66.5 49.1 35.6 43.0
ILLUME [56] $7$B 224 88.5 1445.3 65.1 72.9 - 38.2 37.0
Janus [59] $1.5$B 384 87.0 1338.0 69.4 63.7 59.1 30.5 34.3
@@
Show-o [62] $1.3$B 256 73.8 948.4 - - 48.7 25.1 -
Show-o [62] $1.3$B 512 80.0 1097.2 - - 58.0 26.7 -
LWM [33] $7$B 256 75.2 - - - 44.8 - 9.6
D-Dit [30] $2.0$B 256 84.0 1124.7 - - 59.2 - -
Emu3333-Chat [58] $8$B 512 85.2 1244.0 58.5 68.2 60.3 31.6 37.2
Chameleon [50] $7$B 256 - - - - - 22.4 8.3
Orthus [23] $7$B 256 79.6 1265.8 - - 52.8 28.2 -
TokenFlow-B [42] $13$B 224 84.0 1353.6 55.3 60.4 59.3 34.2 22.4
TokenFlow-L [42] $13$B 256 85.0 1365.4 60.3 62.6 60.3 34.4 27.7
SynerGen-VL [28] $2.4$B 512 85.3 1381.0 53.7 62.0 59.7 34.2 34.5
VILA-U [61] $7$B 256 83.9 1336.2 - 56.3 48.3 - 27.7
VILA-U [61] $7$B 384 85.8 1401.8 - 59.0 60.8 - 33.5
TokLIP (Ours) $7$B 384 84.1 1488.4 67.6 70.4 59.5 43.1 29.8
@@
To assess the effectiveness of different tokenizer architectures integrated into MLLMs, we compare TokLIP with VILA-U [61] and VQGAN [47] on multimodal comprehension tasks. Specifically, we use the LLaVA-v1.5 [35] framework to evaluate the understanding performance of VQGAN and TokLIP, with VILA-U results reported from the original paper. We train on the LLaVA-Pretrain-558K dataset for MLP pretraining and LLaVA-v1.5-mix-665K for instruction tuning. The comprehensive evaluations are presented in Table 1. Our results demonstrate that TokLIP consistently outperforms VQGAN across all metrics and on nearly all benchmarks. For 256 input resolution, TokLIP achieves a clear performance advantage over both VILA-U and VQGAN. Notably, VILA-U is trained on 7M vision-language pairs, significantly more than the LLaVA-v1.5 dataset. Despite these variations in training data size, input resolution changes have a negligible impact on VQ-based tokenizers’ performance, pointing to the inherent limitations in scaling their comprehension capacity. These findings further support our hypothesis that: (1) discrete tokenizers struggle with comprehension tasks due to their lack of semantic depth, and (2) discrete semantic tokenizers face performance degradation due to conflicting training objectives. Overall, our results highlight TokLIP’s superior ability to capture semantic-level understanding, making it a more effective solution for multimodal comprehension tasks involving discrete inputs.
@@ 4.3 Multimodal Comprehension
We present a comprehensive evaluation of TokLIP in comparison to state-of-the-art models in both understanding-only MLLMs and unified understanding-and-generation MLLMs. Illustrated in Table 3, TokLIP, as a discrete tokenizer, achieves competitive performance when compared to existing methods that rely on continuous inputs from CLIP vision encoders. For instance, TokLIP performs on par with LLaVA-v1.5 across most benchmarks, while outperforming it on the SEED Bench, MMB, and MMMU datasets. These results demonstrate that TokLIP effectively learns semantic features from discrete visual tokens, significantly enhancing multimodal understanding capabilities. By leveraging the pre-trained CLIP model, TokLIP successfully incorporates high-level semantic information, bridging the gap between visual tokens and CLIP-level semantics.
In the context of unified models with discrete inputs, which represents the most comparable setting, TokLIP substantially outperforms Chameleon and Show-o across all evaluated benchmarks. Moreover, TokLIP surpasses Emu3-chat on most benchmarks, despite utilizing much less training data and operating at a lower input resolution (384 vs. 512). This reinforces the effectiveness of TokLIP in successfully semanticizing VQ codes with CLIP-level semantics, showing that high-quality semantic understanding can be achieved with fewer resources. When compared to VILA-U, TokLIP also delivers superior performance, further supporting our hypothesis that transitioning from discrete to continuous semantics offers a more effective approach, as discussed in Section 3.3. These empirical results highlight TokLIP’s potential as a foundational tokenizer for unified understanding and generation tasks, paving the way for more advanced multimodal models.
@@
Approaches [58, 62] that directly optimize discrete VQ-based tokenizers typically rely on large-scale datasets. For example, SynerGen-VL [28] is trained on 170 million samples with task-specific prompts for visual understanding, which is often unaffordable for most researchers. Similarly, discrete semantic models [61, 42, 71] like VILA-U require large-scale training sets to discretize high-level semantics effectively. In contrast, we emphasize that TokLIP provides an efficient solution for tokenizer training, achieving strong performance with just 80M image-text pairs and fewer resources for visual understanding. This efficiency is due to our pipeline, which bridges the gap from discrete tokens to continuous representations, enabling effective learning with less data and computational overhead.
@@ 4.4 Visual Generation
We follow the training pipeline of LlamaGen [47] to evaluate the generation capacity on ImageNet benchmarks. Specifically, we replace the text-to-image VQGAN tokenizer with our TokLIP tokenizer to train both GPT-Base and GPT-Large models for different epochs, while keeping the training settings and hyperparameters consistent. We compare the discrete VQGAN tokenizer with TokLIP in Table 4. All experiments are conducted with the same sampling configuration: cfg = 2.5, top-k = 0 (all), top-p = 1.0, and temperature = 1.0. Our results show that all models using the TokLIP tokenizer achieve a lower FID than those using the discrete VQGAN tokenizer, demonstrating that the incorporation of high-level semantics enhances image generation. The improvement is less significant at the 256 input image size compared to the 384 resolution, which we attribute to the fact that semantics are more beneficial when dealing with longer image tokens.

### v2

@@ 4.1 Experimental Setup
The TokLIP architecture integrates a pre-trained text-to-image VQGAN tokenizer with a causal token encoder based on the ViT-So400M architecture. We implement three variants of TokLIP with different input resolutions: TokLIP-B (256×256), TokLIP-L (384×384), and TokLIP-XL (512×512). TokLIP-B and TokLIP-L use the VQGAN from LlamaGen (Sun et al., 2024a), while TokLIP-XL adopts IBQ (Shi et al., 2024a) with a 260K-sized codebook. All models operate with a downsampling ratio of 16. The token encoder is initialized from SigLIP2 (Tschannen et al., 2025), providing a strong foundation for semantic understanding. A two-layer MLP bridges the VQGAN output and the token encoder, projecting the low-dimensional VQGAN features into the input space of the CLIP-based encoder. For training, TokLIP-L and TokLIP-XL are trained on 80/70 million image-text pairs from CapsFusion  (Yu et al., 2024), CC12M (Changpinyo et al., 2021), and LAION-high-resolution (Schuhmann et al., 2022). TokLIP-B further includes a subset of LAION400M (Schuhmann et al., 2021), resulting in a combined dataset of 125 million samples. Low-resolution images are filtered to ensure effective model learning.
For multimodal comprehension, we utilize TokLIP as the input encoder and pass the output concatenated features through a 2-layer MLP to Qwen2.5-7B-Instruct (Yang et al., 2024). Our training process consists of two stages. In Stage 1, we jointly train the MLP and the LLM on 4.5 million images with detailed captions for multimodal alignment, using data sourced from LLaVA-OneVision (Li et al., 2024a). In stage 2, we perform instruction tuning using the LLaVA-OneVision along with LLaVA-Next (Liu et al., 2024c) datasets. For visual generation, we follow LlamaGen (Sun et al., 2024a), training class-to-image generation models on ImageNet. We replace LlamaGen’s text-to-image VQGAN tokenizer with TokLIP.
@@
We evaluate the understanding capacity of our TokLIP on ImageNet classification (Deng et al., 2009) and MSCOCO 5K retrieval (Chen et al., 2015) tasks. For multimodal comprehension, we rigorously evaluate our model on several zero-shot vision-language benchmarks, which include POPE Li et al. (2023b), MME-P (MME-Perception) (Fu et al., 2023), MMB (MMBench) (Liu et al., 2024d), SEED (SEED-Bench Img) (Li et al., 2023a), GQA (Hudson and Manning, 2019), MMMU (Yue et al., 2024), and MM-Vet (Yu et al., 2023b), covering a diverse range of tasks that test both visual and linguistic understanding in a variety of real-world settings. For visual generation, we mainly report Fréchet inception distance (FID) on the ImageNet 256x256 benchmark.
Table 2: Comparative analysis of tokenizers on multimodal comprehension tasks. We adopt the LLaVA-v1.5 (Liu et al., 2024b) training framework and combine tokenizers with Qwen2.5-7B-Instruct (Yang et al., 2024), training on LLaVA-v1.5 data. denotes the results are reported from the original paper.
@@ 4.2 Comparison with Other Tokenizers
We compare TokLIP with state-of-the-art tokenizers (Wu et al., 2024b; Zhao et al., 2025) and CLIP models (Cherti et al., 2023; Radford et al., 2021; Zhai et al., 2023) across both image classification and cross-modal retrieval tasks, shown in Table 1. It can be observed that TokLIP outperforms several SOTA CLIP-based encoders. For instance, TokLIP achieves an impressive 76.4% accuracy on ImageNet with a 256 input resolution, surpassing MetaCLIP (Xu et al., 2023) and EVA02 (Fang et al., 2024). These results demonstrate that TokLIP effectively inherits semantic understanding from the pre-trained teacher and demonstrates competitive performance on downstream tasks. Additionally, at both 256x256 and 384x384 input resolutions, TokLIP outperforms VILA-U (Wu et al., 2024b) and QLIP (Zhao et al., 2025), particularly excelling in the retrieval task. TokLIP achieves 68.00% TR@1 and 52.87% IR@1, which are 7.14% and 9.87% higher than QLIP, respectively, at a 384 resolution. This suggests that TokLIP captures more semantic information compared to methods that discretize high-level image features. Notably, VILA-U is trained on COYO-700M (Byeon et al., 2022), and QLIP uses DataComp-1B (Gadre et al., 2023), both of which involve significantly more resource-intensive training processes than ours. Additionally, TokLIP-XL outperforms SigLIP-B/16, which was trained on the 10B-scale WEBIL dataset, demonstrating that high-resolution inputs combined with a stronger VQGAN tokenizer can further boost comprehension performance. Future work will focus on enhancing TokLIP’s performance through the utilization of higher-quality training data (Fang et al., 2023; Xu et al., 2023).
@@
InstructBLIP Dai et al. (2023) Vicuna-7B 224 - - 36.0 53.4 49.2 30.6 33.8
InstructBLIP Dai et al. (2023) Vicuna-13B 224 78.9 1212.8 - - 49.5 - -
Qwen-VL-Chat Bai et al. (2023) Qwen-7B 448 - 1487.5 60.6 58.2 57.5 35.9 45.9
LLaVA-v1.51.5 Liu et al. (2024b) Vicuna1.5-7B 336 85.9 1510.7 64.3 58.6 62.0 35.4 54.8
@@
DreamLLM Dong et al. (2023) Vicuna1.5-7B 224 - - 58.2 - - - -
LaVIT Jin et al. (2023) LLaMA-7B 224 - - 58.0 - 46.8 - -
MetaMorph Tong et al. (2024) LLaMA3-8B 384 - - 75.2 71.8 - - -
NExT-GPT Wu et al. (2023) Vicuna-7B 224 - - - 57.5 - - -
SEED-X Ge et al. (2024) LLaMA2-chat-13B 448 84.1 1457.0 70.1 66.5 49.1 35.6 -
ILLUME Wang et al. (2024a) Vicuna1.5-7B 224 88.5 1445.3 65.1 72.9 - 38.2 -
Janus Wu et al. (2024a) DeepSeek-LLM-1.5B 384 87.0 1338.0 69.4 63.7 59.1 30.5 -
Janus-Pro-1B Chen et al. (2025c) DeepSeek-LLM-1.5B 384 86.2 1444.0 75.5 68.3 59.3 36.3 -
MAR Wu et al. (2025) Qwen-2.5-1.5B 512 87.6 1155.0 65.5 67.1 58.9 38.9 -
@@
Show-o Xie et al. (2024a) Phi-1.5-1.3B 256 73.8 948.4 - - 48.7 25.1 -
Show-o Xie et al. (2024a) Phi-1.5-1.3B 512 80.0 1097.2 - - 58.0 26.7 -
LWM Liu et al. (2024a) LLaMA2-7B 256 75.2 - - - 44.8 - -
D-Dit Li et al. (2024c) 22B diffusion 512 84.0 1124.7 - - 59.2 - -
Emu33-Chat Wang et al. (2024b) 88B from scratch 512 85.2 1244.0 58.5 68.2 60.3 31.6 70.0
Chameleon Team (2024) 77B from scratch 256 - - - - - 22.4 46.0
Orthus Kou et al. (2024) Chameleon-7B 256 79.6 1265.8 - - 52.8 28.2 -
VILA-U Wu et al. (2024b) LLaMA2-7B 256 83.9 1336.2 - 56.3 48.3 - -
VILA-U Wu et al. (2024b) LLaMA2-7B 384 85.8 1401.8 - 59.0 60.8 - -
UniTok Ma et al. (2025a) LLaMA2-7B 256 83.2 1448.0 - 61.1 - - -
UniToken Jiao et al. (2025) Chameleon-7B 384 - - 71.1 69.9 - 32.8 68.7
SemHiTok Chen et al. (2025d) Vicuna1.5-7B 384 84.2 1400.6 60.3 62.4 61.0 35.2 -
SynerGen-VL Li et al. (2024b) InternLM2-2.4B 512 85.3 1381.0 53.7 62.0 59.7 34.2 60.8
MUSE-VL Xie et al. (2024b) Qwen2.5-7B 256 - - 72.1 69.1 - 39.7 69.8
TokenFlow-L Qu et al. (2024) Vicuna1.5-13B 256 85.0 1365.4 60.3 62.6 60.3 34.4 56.6
TokenFlow-XL Qu et al. (2024) Vicuna1.5-13B 384 86.8 1545.9 68.9 68.7 62.7 38.7 66.7
TokLIP-L (Ours) Qwen2.5-7B 384 84.9 1496.6 76.9 71.5 57.0 47.1 76.8
TokLIP-XL (Ours) Qwen2.5-7B 512 85.2 1586.8 77.4 72.1 57.3 47.1 77.7
@@
To assess the effectiveness of different tokenizer architectures integrated into MLLMs, we compare TokLIP with VILA-U (Wu et al., 2024b) and VQGAN (Sun et al., 2024a) on multimodal comprehension tasks. Specifically, we use the LLaVA-v1.5 (Liu et al., 2024b) framework to evaluate the understanding performance of VQGAN and TokLIP, with VILA-U results reported from the original paper. We train on the LLaVA-Pretrain-558K dataset for MLP pretraining and LLaVA-v1.5-mix-665K for instruction tuning. The comprehensive evaluations are presented in Table 2. Our results demonstrate that TokLIP consistently outperforms VQGAN across all metrics and on nearly all benchmarks. For 256 input resolution, TokLIP achieves a clear performance advantage over both VILA-U and VQGAN. Notably, VILA-U is trained on 7M vision-language pairs, significantly more than the LLaVA-v1.5 dataset. Despite these variations in training data size, input resolution changes have a negligible impact on VQ-based tokenizers’ performance, pointing to the inherent limitations in scaling their comprehension capacity. These findings further support our hypothesis that: (1) discrete tokenizers struggle with comprehension tasks due to their lack of semantic depth, and (2) discrete semantic tokenizers face performance degradation due to conflicting training objectives. Overall, our results highlight TokLIP’s superior ability to capture semantic-level understanding, making it a more effective solution for multimodal comprehension tasks involving discrete inputs.
@@ 4.3 Multimodal Comprehension
We present a comprehensive evaluation of TokLIP in comparison to state-of-the-art models in both understanding-only MLLMs and unified understanding-and-generation MLLMs. Illustrated in Table 3, TokLIP, as a discrete tokenizer, achieves competitive performance when compared to existing methods that rely on continuous inputs from CLIP vision encoders. For instance, TokLIP-L performs on par with LLaVA-v1.5 across MME and POPE, while outperforming it on the SEED Bench, MMB, MMMU, and AI2D datasets. These results demonstrate that TokLIP effectively learns semantic features from discrete visual tokens, significantly enhancing multimodal understanding capabilities. By leveraging the pre-trained CLIP model, TokLIP successfully incorporates high-level semantic information, bridging the gap between visual tokens and CLIP-level semantics.
In the context of unified models with discrete inputs, which represents the most comparable setting, TokLIP-L substantially outperforms Chameleon and Show-o across all evaluated benchmarks. Moreover, TokLIP-L surpasses Emu3-chat on most benchmarks, despite utilizing much less training data and operating at a lower input resolution (384 vs. 512). This reinforces the effectiveness of TokLIP in successfully semanticizing VQ codes with CLIP-level semantics, showing that high-quality semantic understanding can be achieved with fewer resources. When compared to VILA-U and TokenFlow, TokLIP-L also delivers superior performance, further supporting our hypothesis that transitioning from discrete to continuous semantics offers a more effective approach, as discussed in Section 3.3. Moreover, TokLIP-XL further boosts comprehension performance across most benchmarks, achieving state-of-the-art results on AI2D, MME, SEED-Bench, and MMB. The sub-optimal performance on GQA may be attributed to the data mixup strategy in LLaVA-OneVision, and we expect further improvements with training on more QA pairs. These empirical results highlight TokLIP’s potential as a foundational tokenizer for unified understanding and generation tasks, paving the way for more advanced multimodal models.
@@
Approaches (Wang et al., 2024b; Xie et al., 2024a) that directly optimize discrete VQ-based tokenizers typically rely on large-scale datasets. For example, SynerGen-VL (Li et al., 2024b) is trained on 170 million samples with task-specific prompts for visual understanding, which is often unaffordable for most researchers. Similarly, discrete semantic models (Wu et al., 2024b; Qu et al., 2024; Zhao et al., 2025) like VILA-U require large-scale training sets to discretize high-level semantics effectively. In contrast, we emphasize that TokLIP provides an efficient solution for tokenizer training, achieving strong performance with just 80M image-text pairs and fewer resources for visual understanding. Note that both TokLIP and VILA-U are initialized from a pre-trained SigLIP variant, so we exclude the SigLIP pretraining data and discuss solely on the tokenizer training stage. This efficiency is due to our pipeline, which bridges the gap from discrete tokens to continuous representations, enabling effective learning with less data and computational overhead.
@@ 4.4 Visual Generation
We follow the training pipeline of LlamaGen (Sun et al., 2024a) to evaluate the generation capacity on ImageNet benchmarks. Specifically, we replace the text-to-image VQGAN tokenizer with our TokLIP tokenizer to train both GPT-Base and GPT-Large models for different epochs, while keeping the training settings and hyperparameters consistent. We compare the discrete VQGAN tokenizer with TokLIP in Table 4. All experiments are conducted with the same sampling configuration: cfg = 2.5, top-k = 0 (all), top-p = 1.0, and temperature = 1.0. Our results show that all models using the TokLIP tokenizer achieve a lower FID than those using the discrete VQGAN tokenizer, demonstrating that the incorporation of high-level semantics enhances image generation. The improvement is less significant at the 256 input image size compared to the 384 resolution, which we attribute to the fact that semantics are more beneficial when dealing with longer image tokens.